OpenStudy (anonymous):

Simplify 4∜48 a. 8∜3 b. 4∜3 c. 12∜4 d. 18∜2

OpenStudy (campbell_st):

well 48 = 16 x 3 and 16 = 2^4 so you can have \[4\times \sqrt[4]{16 \times 3} = 4 \times \sqrt[4]{16} \times \sqrt[4]{3} = 4 \times 2 \times \sqrt[4]{3}\]

OpenStudy (anonymous):

Okay, so it's a. So the 4 over the radical doesn't really do anything?

OpenStudy (campbell_st):

it just means the 4th root... instead of the cube or square root

OpenStudy (anonymous):

Oh, Okaay, thank you! (: